OpenStudy (abhisar):

A small sphere is suspended by a string from the ceiling of a car. If the car begins to move with a constant acceleration a, prove that the inclination of the string to the vertical is \(\sf tan^{-1} (a/g)\)

OpenStudy (anonymous):

  \[T \sin \theta = m a\] \[m = \frac{ T \sin \theta }{a }\] \[T \cos \theta = m g\] \[m = \frac{ T \cos \theta }{ g }\] thus \[\frac{ T \cos \theta }{ g } = \frac{ T \sin \theta }{ a }\] \[\tan \theta = \frac{ a }{ g }\] \[\theta = \tan^{-1} (\frac{ a }{ g })\] Q.E.D.